# CARLA: Self-supervised Contrastive Representation Learning for Time Series Anomaly Detection

### v3

@@ 4.5 Ablation Study
Additionally, for a deeper understanding of the individual contributions of the stages and different components in our proposed model, we conduct an ablation study. Our analysis focused on the following aspects: (i) Effectiveness of CARLA’s two stages, (ii) Positive pair selection strategy (iii) Effectiveness of Different Anomaly Types (iv) Effectiveness of the loss components.
@@ 4.5.1 Effectiveness of CARLA’s Stages
In this section, we delve into an in-depth evaluation of the stages of our proposed model, utilising the M-6 time series derived from the MSL dataset. Our primary objective is to develop a comprehensive grasp of the patterns inherent in the features extracted and learned by our model throughout its stages. For this purpose, we employ t-distributed Stochastic Neighbor Embedding (t-SNE) to visualise the output of the Pretext Stage and the Self-supervised Classification stage. These are graphically represented in Figures 5(a) and 5(b) correspondingly. From these graphical representations, it is palpable that the second stage significantly enhances the discrimination capacity between normal and anomalous samples. It is important to highlight that the first stage reveals some semblance of anomalous samples to normal ones, which might be due to the existence of anomalies close to normal boundaries. However, the Self-supervised Classification stage efficiently counteracts this ambiguity by effectively segregating these instances, thereby simplifying subsequent classification tasks.
Figure 6: Comparative analysis of the model’s stages utilising t-SNE and anomaly score distributions on M-6 dataset in MSL. (a) t-SNE after the Pretext stage, indicating the initial representation of normal and anomalous samples. (b) t-SNE after the Self-supervised Classification stage, demonstrating enhanced segregation of normal and anomalous instances. (c) Anomaly score distribution in the Pretext Stage based on Euclidean distances between test and training samples. (d) The distribution of anomaly scores following the application of the Self-supervised Classification Stage, evidencing an improved representation.
@@
In the Pretext Stage, the anomaly score is computed as the minimum distance between the test sample and all original training samples in the representation space. This approach aims to identify the closest match among the training samples for the given test sample. A smaller distance implies a greater probability that the test sample belongs to the normal data distribution.
@@
As we can observe from the figures 5(a) and 5(b), the self-supervised classification Stage has resulted in a significant improvement in the separation of normal and anomalous windows. The distribution of anomaly scores in Figure 5(d) is more clearly separated than in Figure 5(c). This indicates that the second stage has produced more consistent representations. Furthermore, the anomaly scores in Figure 5(d) are relatively closer to 0 or 1, which indicates an improvement in discrimination between normal and anomalous windows in the second stage of CARLA.
@@ 4.5.2 Positive Pair Selection
Our experiment demonstrated that selecting positive pairs using a random temporal neighbour is a more effective approach than weak augmentation with noise. Since anomalies occur rarely, choosing a positive pair that is in the target window’s temporal proximity is a better strategy in our model.
@@ 4.5.3 Effectiveness of Different Anomaly Types
Based on the results of the ablation study on the MSL dataset, CARLA’s performance is evaluated by systematically removing different types of anomalies during the Pretext Stage. The evaluation metrics used are precision, recall, F1 score, and AU-PR. The results are sorted based on F1 from the least significant to the most significant types of anomalies in Table 5. This result represents the overall performance of the CARLA model when using all types of anomalies during the anomaly injection process. It serves as a baseline for comparison with the subsequent results.
@@
Eliminating trend anomalies has a significant negative impact on CARLA’s precision. However, the recall remains relatively high, indicating that trend anomalies are important for maintaining a higher precision level. Similar to the previous case, removing contextual anomalies leads to a slight decrease in precision but an improvement in recall. The exclusion of shapelet anomalies leads to a moderate decrease in precision, while the recall remains relatively high. This suggests that shapelet anomalies contribute to the model’s precision but are not as crucial for capturing anomalies in general.
By removing global anomalies, CARLA’s precision slightly decreases, indicating that it becomes less accurate in identifying true anomalies. However, the recall improves, suggesting that the model becomes more sensitive in detecting anomalies overall. Removing seasonal anomalies has a noticeable effect on the model’s precision, which drops significantly. The consistently high recall suggests the model effectively captures many anomalies without relying on seasonal patterns.
Overall, the analysis of the results suggests that global and seasonal anomalies play a relatively more significant role in the CARLA model’s performance, while contextual, shapelet, and trend anomalies have varying degrees of impact. These findings provide insights into the model’s sensitivity to different types of anomalies and can guide further improvements in anomaly injection strategies within CARLA’s framework.
@@ 4.6 Parameter Sensitivity
We also examine the sensitivity of CARLA’s parameters. We investigate and analyze the effect of the window size, number of classes, number of NN/FN, and coefficient of entropy.
@@ 4.6.1 Effect of Window Size
WS
Prec
Rec
F1
AUPR
Prec
Rec
F1
AUPR
Prec
Rec
F1
AUPR
50
0.2755
0.7702
0.4058
0.3173
0.4569
0.4885
0.4722
0.4636
0.3046
0.8800
0.4526
0.5533
100
0.3233
0.7666
0.4548
0.4155
0.2584
0.5938
0.3598
0.4136
0.3709
0.9713
0.5368
0.5331
150
0.3606
0.7823
0.4936
0.4595
0.3599
0.6592
0.4655
0.4661
0.5720
0.9578
0.7162
0.6685
200
0.3891
0.7959
0.5227
0.5009
0.4276
0.6362
0.5114
0.5070
0.5747
0.9755
0.7233
0.6450
250
0.4380
0.8037
0.5412
0.5188
0.4224
0.6596
0.5150
0.4850
0.7129
0.9781
0.8247
0.7164
Based on the analysis of both the F1 score and AU-PR, it can be concluded that window size 200 consistently outperforms other window sizes on the MSL, SMD, and Yahoo datasets overall. This window size strikes a balance between precision and recall, effectively capturing anomalies while maintaining a high discrimination ability. Therefore, window size 200 is selected as the best choice for conducting time series analysis in all experiments. Additionally, Figure 7 shows that our model’s performance is stable in window sizes between 150 and 250.
@@
4.6.2 Effect of Number of Classes
The provided results showcase the performance of the model with varying numbers of classes on the MSL dataset. Based on the illustration of both the F1 score and AU-PR in Figure 7(a), the model performs relatively well across different numbers of classes. While CARLA’s performance is pretty stable across all number of classes denoted on the x-axis of the plot in Figure 7(a), it can be concluded that using 10 classes as the number of class parameters yields the highest performance. It strikes a balance between capturing anomalies and minimising false positives while effectively discriminating between anomalies and normal data points.
The results suggest that increasing the number of classes beyond 10 does not significantly improve the model’s performance. When the number of classes increases, normal representations divide into the different classes and, in CARLA, are detected as anomalies (lower probability of belonging to the major class). However, using 2 classes leads to a lower performance compared to the selected parameter of 10 classes, implying that a more fine-grained classification with 10 classes provides the assumption that various anomalies are spread in the representation space.
@@ 4.6.3 Effect of Number of neighbours in $\mathcal{N}$ and $\mathcal{F}$
Based on the analysis of both the F1 score and AU-PR, we can see that while CARLA’s performance is pretty stable across the different numbers of neighbours in $\mathcal{N}$ and $\mathcal{F}$ denoted on the x-axis of the plot in Figure 7(b), it can be concluded that employing 5 nearest/furthest neighbours outperforms the other options. This number of parameters strikes a favourable balance between precision and recall, allowing for effective anomaly detection while maintaining a high discrimination ability. Therefore, selecting 5 for the number of neighbours is deemed the optimal choice for this experiment, as it consistently achieves higher F1 scores and AU-PR values.
4.6.4 Effect of Coefficient of Entropy
@@
Overall, the analysis indicates that coefficients of 2, 5, and 10 are viable for the entropy component in the loss function, but a coefficient of 5 is chosen as optimal due to its superior F1 score and AU-PR.
@@ 5 Conclusion
In conclusion, our innovative end-to-end self-supervised framework, CARLA, utilises contrastive representation learning with anomaly injection to generate robust representations for time series windows and classify anomalous windows. The use of semantically meaningful nearest and furthest neighbours as a prior allows us to capture underlying patterns in time series data and learn a representation that is well-aligned with these patterns, thereby enhancing the accuracy of anomaly detection. Additionally, the experimental evaluation shows promising results in detecting anomalies in the seven most commonly used real-world time series datasets, demonstrating the effectiveness of CARLA in this domain.
This research has demonstrated the potential of contrastive learning combined with synthetic anomaly injection to overcome the limitations of the lack of labelled data in time series anomaly detection. To enhance our model’s performance, future research will focus on refining positive and negative sample selection strategies. Moreover, we believe that there is much scope to improve performance by broadening the forms of injected anomalies used in our framework. We further hypothesise that varying the severity of injected anomalies may improve the model’s ability to detect anomalies in representation space. We commend the investigation of these intriguing prospects in future research.
Additionally, considering that some datasets with a significant number of abnormal samples, resulting in more anomalous windows during the Pretext Stage, it becomes critical to revisit the impact of selecting positives and negatives for abnormal anchor samples under unsupervised settings. This reconsideration is crucial for enhancing the model’s effectiveness in time series anomaly detection.
In conclusion, this paper’s results, based on appropriate metrics, contribute significantly to further investigations and advancements in time series anomaly detection, particularly on datasets like WADI and KPI. By employing the correct evaluation metrics, such as the conventional F1 score without point adjustment and AU-PR, this research provides a solid foundation for assessing the performance of future models in this field. The findings presented herein enable researchers to benchmark their models effectively and compare their results against established standards. Reliable evaluation metrics and validated results serve as valuable resources, driving progress and innovation in time series anomaly detection research.
@@ Appendix A List of Symbols and Definitions
$\mathcal{D}$
Sequential time series windows where $\mathcal{D}=\{w_{1},w_{2},..,w_{m}\}$
$w_{i}$
$i^{th}$ window in $\mathcal{D}$ where $w_{i}=\{x_{1},x_{2},..,x_{WS}\}$
$WS$
Time series window size
$Dim$
Number of dimensions in time series $\mathcal{D}$
$m$
Number of windows
$Q$
Number of nearest/furthest neighbours
$a$
Anchor sample (i.e. the original time series window)
$y$
Number of temporally proximate windows
$p$
Positive sample (i.e. random window selected from $y$ temporally proximate previous windows)
$n$
Negative sample (i.e. anomaly injected window)
$(a,p,n)$
Triplet in Pretext Stage
$\mathcal{T}$
All triplets set in Pretext Stage
$\alpha$
Margin in pretext loss
$\phi_{p}$
Pretext trained model
$\mathcal{B}$
All neighbours set
$\mathcal{N}$
$Q$-Nearest neighbours set
$\mathcal{F}$
$Q$-Furthest neighbours set
$\phi_{s}$
Self-supervised trained model
$C_{m}$
Majority class label
$\beta$
Entropy loss weight
Appendix B Point Adjustment (PA)
A notable issue with many existing Time Series Anomaly Detection (TSAD) methods is their reliance on a specific evaluation approach called point adjustment (PA) strategy (Xu et al., 2018) and utilised in TSAD literature such as Su et al. (2019). PA operates under the following principle: if a given segment is detected as anomalous, all previous segments in that anomaly subsequence are considered as anomalies (see Figure 9). PA was introduced under the assumption that a solitary alert during an anomaly phase is adequate to recover the system. It has evolved into a fundamental stage in the assessment of TSAD, with certain studies exclusively documenting $F1_{PA}$ without the conventional F1 score Kim et al. (2022). A higher $F1_{PA}$ is often interpreted as indicating better detection capability.
Figure 9: Impact of point adjustment (PA) on anomaly detection in time series. The figure represents a toy example using a sequence of 12 samples. Presents the anomaly scores for each sample within the sequence. The second row depicts the outcomes of anomaly detection using a threshold of 0.5 applied to all samples’ anomaly scores. The third row provides the actual labels (ground truth) for the samples within the sequence. The fourth and final row presents the adjusted anomaly detection results using point adjustment with the ground truth. Evaluation metrics are calculated for both unadjusted and adjusted anomaly detection.
In Figure 9, for the unadjusted detection, we have True Positives (TP): 2, True Negatives (TN): 3, False Positives (FP): 1, False Negatives (FN): 6, yielding a Recall of 0.25, Precision of 0.67, and F1 score of 0.36. In contrast, the adjusted detection (in bold for emphasis) resulted in TP: 7, TN: 3, FP: 1, and FN: 1, giving a significantly improved Recall of 0.875, Precision of 0.875, and F1 score of 0.875. This illustrates the substantial difference in evaluation results due to the application of PA in the anomaly detection process. It is evident that the point adjustment greatly enhances the performance of anomaly detection.
Table 9: Precision ($Prec_{PA}$), recall ($Rec_{PA}$), and $F1_{PA}$ results with point adjustment (PA) for various models on multivariate time series datasets.
@@ Appendix B Point Adjustment (PA)
Table 10: Precision ($Prec_{PA}$), recall ($Rec_{PA}$), and $F1_{PA}$ results with point adjustment (PA) for various models on univariate time series datasets.
@@
However, PA carries a notable peril of inflating the performance of the model. In most of the TSAD models, an anomaly score is produced to indicate the degree of abnormality, and the model detects a given point as an anomaly when the score surpasses a predefined threshold. Incorporating PA, the predictions based on the model that randomly generated anomaly scores uniformly outperform all state-of-the-art models, as depicted in Tables 9 and 10. It becomes challenging to assert that a model with a superior $F1_{PA}$ performs better than a model with a lower $F1_{PA}$ when random anomaly scores have the potential to result in a high $F1_{PA}$ comparable to that of a proficient detection model. Experimental results in Tables 9 and 10 demonstrate that the random anomaly score model can outperform all state-of-the-art models, highlighting the limitations of relying solely on $F1_{PA}$ for model evaluation.

### v4

@@ 4.5 Ablation Study
For a deeper understanding of the contributions of different stages and components in our proposed model, we conduct an ablation study. Our analysis focused on: (i) Effectiveness of CARLA’s two stages, (ii) Positive pair selection strategy (iii) Effectiveness of Different Anomaly Types (iv) Effectiveness of the loss components.
@@ 4.5.1 Effectiveness of CARLA’s Stages
In this section, we delve into an in-depth evaluation of the stages of our proposed model, utilising the M-6 time series derived from the MSL dataset. Our primary objective is to develop a comprehensive grasp of the patterns inherent in the features extracted and learned by our model throughout its stages. For this purpose, we employ t-distributed Stochastic Neighbor Embedding (t-SNE) to visualise the output of the Pretext Stage and the Self-supervised Classification stage. These are graphically represented in Figures 6(a) and 6(b) correspondingly. From these graphical representations, it is palpable that the second stage significantly enhances the discrimination capacity between normal and anomalous samples. It is important to highlight that the first stage reveals some semblance of anomalous samples to normal ones, which might be due to the existence of anomalies close to normal boundaries. However, the Self-supervised Classification stage efficiently counteracts this ambiguity by effectively segregating these instances, thereby simplifying subsequent classification tasks.
Figure 6: Comparative analysis of the model’s stages utilising t-SNE and anomaly score distributions on M-6 dataset in MSL. (6(a)) t-SNE after the Pretext stage, indicating the initial representation of normal and anomalous samples. (6(b)) t-SNE after the Self-supervised Classification stage, demonstrating enhanced segregation of normal and anomalous instances. (6(c)) Anomaly score distribution in the Pretext Stage based on Euclidean distances between test and training samples. (6(d)) The distribution of anomaly scores following the application of the Self-supervised Classification Stage evidencing an improved representation.
@@
In the Pretext Stage, the anomaly score is computed as the minimum distance between the test sample and all original training samples in the representation space, aiming to identify the closest match among the training samples for the given test sample. A smaller distance implies a greater probability that the test sample belongs to the normal data distribution.
@@
As we can observe from the figures 6(a) and 6(b), the self-supervised classification Stage has resulted in a significant improvement in the separation of normal and anomalous windows. The distribution of anomaly scores in Figure 6(d) is more clearly separated than in Figure 6(c). This indicates that the second stage has produced more consistent representations. Furthermore, the anomaly scores in Figure 6(d) are relatively closer to 0 or 1, indicating improved discrimination between normal and anomalous windows in the second stage of CARLA.
@@ 4.5.2 Positive Pair Selection
Our experiment demonstrated that selecting positive pairs using a random temporal neighbour is a more effective approach than weak augmentation with noise. Since anomalies occur rarely, choosing a positive pair that is in the target window’s temporal proximity is a better strategy.
@@ 4.5.3 Effectiveness of Different Anomaly Types
CARLA’s performance is evaluated on the MSL dataset by systematically removing different types of anomalies during the Pretext Stage. The evaluation metrics used are precision, recall, F1 score, and AU-PR. The results are sorted based on F1 from the least significant to the most significant types of anomalies in Table 5. This result represents the overall performance of the CARLA model when using all types of anomalies during the anomaly injection process. It serves as a baseline for comparison with the subsequent results.
@@
Eliminating trend anomalies has a significant negative impact on CARLA’s precision. However, the recall remains relatively high, indicating that trend anomalies are important for maintaining a higher precision level. Similar to the previous case, removing contextual anomalies leads to a slight decrease in precision but an improvement in recall. The exclusion of shapelet anomalies leads to a moderate decrease in precision, while the recall remains relatively high. This suggests that shapelet anomalies contribute to the model’s precision but are not as crucial for capturing anomalies in general. By removing global anomalies, CARLA’s precision slightly decreases, indicating that it becomes less accurate in identifying true anomalies. However, the recall improves, suggesting that the model becomes more sensitive in detecting anomalies overall. Removing seasonal anomalies significantly drops precision, though the high recall indicates effective anomaly detection without relying on seasonal patterns. Overall, the analysis of the results suggests that global and seasonal anomalies play a relatively more significant role in the CARLA model’s performance, while contextual, shapelet, and trend anomalies have varying degrees of impact. These findings provide insights into the model’s sensitivity to different types of anomalies and can guide further improvements in anomaly injection strategies within CARLA’s framework.
@@ 4.6 Parameter Sensitivity
We examine the sensitivity of CARLA’s parameters, investigating and analysing the effect of the window size, number of classes, number of NN/FN, and entropy coefficient.
@@ 4.6.1 Effect of Window Size
WS Prec Rec F1 AUPR Prec Rec F1 AUPR Prec Rec F1 AUPR
50 0.2755 0.7702 0.4058 0.3173 0.4569 0.4885 0.4722 0.4636 0.3046 0.8800 0.4526 0.5533
100 0.3233 0.7666 0.4548 0.4155 0.2584 0.5938 0.3598 0.4136 0.3709 0.9713 0.5368 0.5331
150 0.3606 0.7823 0.4936 0.4595 0.3599 0.6592 0.4655 0.4661 0.5720 0.9578 0.7162 0.6685
200 0.3891 0.7959 0.5227 0.5009 0.4276 0.6362 0.5114 0.5070 0.5747 0.9755 0.7233 0.6450
250 0.4380 0.8037 0.5412 0.5188 0.4224 0.6596 0.5150 0.4850 0.7129 0.9781 0.8247 0.7164
Based on the analysis of both the F1 score and AU-PR, it can be concluded that window size 200 consistently outperforms other window sizes on the MSL, SMD, and Yahoo datasets overall. This window size strikes a balance between precision and recall, effectively capturing anomalies while maintaining a high discrimination ability. Therefore, window size 200 is selected as the best choice in all experiments. Additionally, Figure 7 shows that our model’s performance is stable in window sizes between 150 and 250.
4.6.2 Effect of Number of Classes
The provided results showcase the performance of the model with varying numbers of classes on the MSL dataset. Based on the illustration of both the F1 score and AU-PR in Figure 8(a), the model performs relatively well across different numbers of classes. While CARLA’s performance is pretty stable across all number of classes denoted on the x-axis of the plot in Figure 8(a), it can be concluded that using 10 classes yields the highest performance. It strikes a balance between capturing anomalies and minimising false positives while effectively discriminating between anomalies and normal samples.
The results suggest that increasing the number of classes beyond 10 does not significantly improve the model’s performance. When the number of classes increases, normal representations divide into the different classes and, in CARLA, are detected as anomalies (lower probability of belonging to the major class). However, using 2 classes leads to a lower performance compared to 10 classes, implying that a more fine-grained classification with 10 classes provides the assumption that various anomalies are spread in the representation space.
@@ 4.6.3 Effect of Number of neighbours in $\mathcal{N}$ and $\mathcal{F}$
Based on the analysis of both the F1 score and AU-PR, we can see that while CARLA’s performance is pretty stable across the different numbers of neighbours in $\mathcal{N}$ and $\mathcal{F}$ denoted on the x-axis of the plot in Figure 8(b), it can be concluded that employing 5 nearest/furthest neighbours outperforms the other options. This number of parameters strikes a favourable balance between precision and recall, allowing for effective anomaly detection while maintaining a high discrimination ability. Therefore, selecting 5 neighbours is deemed the optimal choice for this experiment, as it consistently achieves higher F1 scores and AU-PR values.
4.6.4 Effect of Entropy Coefficient
@@
Overall, the analysis indicates that coefficients of 2, 5, and 10 are viable for the entropy component in the loss function, with 5 being optimal due to its superior F1 score and AU-PR.
@@ 5 Conclusion
Our innovative end-to-end self-supervised framework, CARLA, utilises contrastive representation learning with anomaly injection to generate robust representations for time series and classify anomalous windows. The use of semantically meaningful nearest and furthest neighbours as a prior allows us to capture underlying patterns in time series and learn a representation that is well-aligned with these patterns, thereby enhancing detection accuracy. Our extensive experimental evaluation of the seven most commonly used time series benchmarks, encompassing 208 datasets with diverse real-world anomalies, shows promising results in detecting anomalies, demonstrating the effectiveness of CARLA in this domain.
This research has demonstrated the potential of contrastive learning combined with synthetic anomaly injection to overcome the limitations of the lack of labelled data in TSAD. To enhance our model’s performance, future research will focus on refining positive and negative sample selection strategies. Moreover, we believe that there is much scope to overcome the limitations of the anomaly injection component by broadening the forms of injected anomalies used in our framework. We further hypothesise that varying the severity of injected anomalies may improve the model’s ability to detect anomalies in representation space. We commend the investigation of these intriguing prospects in future research.
Additionally, considering that some datasets have a significant number of abnormal samples, resulting in more anomalous windows during the Pretext Stage, it becomes critical to revisit the impact of selecting positives and negatives for abnormal anchors in unsupervised settings. This reconsideration is crucial for enhancing the model’s effectiveness in TSAD.
Our results, based on appropriate metrics, contribute significantly to further investigations and advancements in TSAD, particularly on datasets like WADI and KPI. By employing the correct evaluation metrics, such as the conventional F1 score without point adjustment and AU-PR, this research provides a solid foundation for assessing the performance of future models in this field. The findings presented herein enable researchers to benchmark their models and compare their results against established standards. Reliable evaluation metrics and validated results serve as valuable resources, driving innovation in TSAD research.
@@ Appendix A List of Symbols and Definitions
$\mathcal{D}$ Sequential time series windows where $\mathcal{D}=\{w_{1},w_{2},..,w_{m}\}$
$w_{i}$ $i^{th}$ window in $\mathcal{D}$ where $w_{i}=\{x_{1},x_{2},..,x_{WS}\}$
$WS$ Time series window size
$Dim$ Number of dimensions in time series $\mathcal{D}$
$m$ Number of windows
$Q$ Number of nearest/furthest neighbours
$a$ Anchor sample (i.e. the original time series window)
$y$ Number of temporally proximate windows
$p$ Positive sample (i.e. random window selected from $y$ temporally proximate previous windows)
$n$ Negative sample (i.e. anomaly injected window)
$(a,p,n)$ Triplet in Pretext Stage
$\mathcal{T}$ All triplets set in Pretext Stage
$\alpha$ Margin in pretext loss
$\phi_{p}$ Pretext trained model
$\mathcal{B}$ All neighbours set
$\mathcal{N}$ $Q$-Nearest neighbours set
$\mathcal{F}$ $Q$-Furthest neighbours set
$\phi_{s}$ Self-supervised trained model
$C_{m}$ Majority class label
$\beta$ Entropy loss weight
Table 9: Precision ($Prec_{PA}$), recall ($Rec_{PA}$), and $F1_{PA}$ results with point adjustment (PA) for various models on multivariate and univariate time series datasets.
(a) Multivariate Time Series Datasets
@@
(b) Univariate Time Series Datasets
@@
Appendix B Point Adjustment (PA)
A notable issue with many existing TSAD methods is their reliance on the point adjustment (PA) strategy [31], as used in TSAD literature like [27]. PA operates on the principle that if a segment is detected as anomalous, all previous segments in that anomaly subsequence are also considered anomalies (see Figure 9). Incorporating PA, the Random model outperform all state-of-the-art models, as depicted in Tables 9(a) and 9(b).
Figure 9: Impact of point adjustment (PA) on anomaly detection in time series. The figure represents a toy example with a sequence of 12 samples, presenting the anomaly scores for each sample. The second row depicts the outcomes of anomaly detection using a threshold of 0.5. The third row provides the actual labels (ground truth). The fourth row presents the adjusted anomaly detection results using point adjustment with the ground truth. Evaluation metrics are calculated for both unadjusted and adjusted anomaly detection.